Suppose we introduce pipelining on this machine. Assume that when


introducing pipelining, the clock skew adds 5ns of overhead to each


execution stage.


i. What is the instruction latency on the pipelined machine?


ii. How much time does it take to execute 100 instructions?

1
Expert's answer
2022-03-18T02:26:30-0400

In the pipelined implementation, the length of the pipe stages must all be the same, i.e., the speed of the slowest stage plus overhead. With 5ns overhead it comes to:

i)The length of pipelined stage = MAX (lengths of unpipelined stages) + overhead = 60 + 5 = 65 ns


ii)Instruction latency = 65 ns Time to execute 100 instructions = 65*6*1 + 65*1*99 = 390 + 6435 = 6825 ns